OpenStudy (anonymous):

Lauren saved $84 more than Fiona. 1/5 of Lauren's savings was the same as 3/8 of Fiona's savings. How much did Lauren and Fiona save altogether?

OpenStudy (anonymous):

Solve the following two simultaneous equations:\[\left\{\text{Lauren} \text{=}\text{Fiona}+84,\frac{1}{5}\text{Lauren}=\frac{3}{8 }\text{Fiona}\right\}\]$276 was the total savings.

OpenStudy (anonymous):

Thank you, but I don't understand how you got that answer.

OpenStudy (anonymous):

Prior posting had an error. Replace Lauren in the second equation with Fiona + 84 and then solve the result for Fiona. Does that help you?

OpenStudy (anonymous):

You can get rid of the fractions by multiplying each side by 40.

OpenStudy (anonymous):

?????? i think the equations that u wrote at first was right

OpenStudy (anonymous):

Thanks, robtobey! I got the answer, thanks again!

OpenStudy (anonymous):

You are welcome. In the last few minutes I put together a Mathematica solution which is attached.